OpenStudy (anonymous):

Solve by the method of your choice? x^2+10x=-24

OpenStudy (anonymous):

factor it

OpenStudy (amistre64):

and KB chooses the OS method ;)

OpenStudy (agentnao):

x^2 + 10x + 24 = 0 (x + 6) (x +4) = 0 X = -6 ,-4

OpenStudy (anonymous):

x^2 + 10x + 24 = 0 (x+6) (x+4) x = -6, -4

OpenStudy (anonymous):

OS method, the best one

OpenStudy (anonymous):

I've literally done dozens of these over the past two days and never have came up with what is correct.

OpenStudy (agentnao):

Try getting a tutor! Maybe they can explain it better then the teacher.

OpenStudy (agentnao):

*than

OpenStudy (anonymous):

I just miss one step or order of operation and screws the whold thing up :(

OpenStudy (anonymous):

Thanks for the help

OpenStudy (agentnao):

Oh, yes :/ PEMDAS is your friend! Remember it and you'll do fine :) You're welcome!